OpenStudy (anonymous):

if 2tanx/1-tan^2x=1, then x can equal

OpenStudy (anonymous):

There can be more than 1 answer a) x=5pi/8 + npi b) x=pi/8 + npi c) x=3pi/8 + npi d) x=7pi/8 + npi

OpenStudy (anonymous):

Please clarify: (2tanx)/1 or 2tan(x/1)

OpenStudy (anonymous):

2tanx/(1-tan^2x)

OpenStudy (anonymous):

(2tanx)(1-tan^2x) =1 ...then x can equal? options above

OpenStudy (anonymous):

See, there is an identity : \[\frac{2tanx}{(1-\tan^2x)} = \tan(2x)\] Using this: \[\tan(2x) = 1\] Or you can do is: \[\tan(2x) = \tan(45)\] Now can you solve further??

OpenStudy (anonymous):

haha i have no clue how to do this

OpenStudy (anonymous):

See, \[\tan(\theta) = \tan(x)\] \[\theta = n \pi + \frac{\pi }{4}\] where n belongs to Integers Set...

OpenStudy (anonymous):

can you tell me the answer for that